How many months will it take Php15,000 to earn Php56.25 at 1.5% simple interest rate?

1
Expert's answer
2021-11-17T17:24:45-0500

"SI=\frac {PRT} {100}"


"56.25=\frac {15,000×1.5×T}{100}"


"5625=22,500T"


"T=\frac {5625}{22500}"


T=0.25


T=3 months